OpenStudy (anonymous):

im not sure about my answer help please - The original function is y = f(x). If the function translates 6 units left and 4 units up, which function represents the translation?

OpenStudy (anonymous):

A. Y=F(X - 6) +4 B. Y=F(X+6)+4 C. Y=F(X-6)-4 D. Y=F(X+4)-6

OpenStudy (twotreats):

I think is B